# First draft: March 2, 2026

### 02 Mar 2026 v1

@@ 1. Introduction
where, $\mathcal{L}$ is the loss function and $\pi(\cdot)$ is the prior. A key implication of loss-based updating is that the scale parameter $\lambda$ is not pinned down by Bayesian conditioning under a likelihood. In standard Bayesian analysis, the likelihood fixes the unit of fit and therefore fixes the update’s scale. Since here, I replace the likelihood with a utility function, I have to choose a learning-rate (or “temperature”) parameter, $\lambda$, and its calibration "moves from a modeling detail to a structural component of principled analysis" (McAllin and Takanashi 2026, p. 2). Rather than treating $\lambda$ as a nuisance parameter, I treat it as the central regularization device, and solve for the optimal value of $\lambda$, $\lambda^{\ast}$, in sample.
@@
I obtain posteriors from US equities over the period 1955-2024. I confirm Lamoureux and Zhang’s finding that characteristic tilts afford large utility gains prior to 2001 for investors with varying tolerance for risk. These gains vanish in the 21st Century. I find that $\lambda^{*}$ varies meaningfully with risk aversion and departs from the quadratic benchmark in ways that are informative about higher-order moments.. The pattern of departure is stable prior to 2001, and changes in ways consistent with structural changes in the data generating process in the 21st Century.
@@
The paper also contributes to the burgeoning literature that represents returns and risks in characteristic space. Like Kelly, Pruitt, and Su’s (2019) Instrumented Principal Components Analysis (IPCA), I shift the investment universe from a set of CUPIPs to a cloud of observations in characteristic space. In contrast to IPCA, which uses this transformation to identify latent statistical factors, I use the geometry of characteristic-managed returns to construct a Gibbs posterior over utility-maximizing portfolios. Kozak, Nagel, and Santosh (2018, 2020) argue that characteristics span the investment opportunity set and estimate the associated low-dimensional structure via shrinkage and low-rank methods. In a complementary way, their work tells us what returns look like in characteristic space and IPCA tells us how risk moves in characteristic space; this paper tells us how to optimally invest in characteristic space under uncertainty.
@@ 3. Bayesian updating
Brandt, Santa-Clara, and Valkanov (2009) stress that an investor can use this algorithm to maximize in-sample utility without taking a stand on the data generating process. By reducing the portfolio problem to that of estimating the $K$-vector $\theta$ of characteristic tilts, thereby avoiding the estimation of moments, or a complex likelihood function, they hope to mitigate the problem of overfitting that plagues traditional portfolio optimization. As described above, as a Bayesian decision maker, I also want to be agnostic about the data generating process—not simply to make the estimation more parsimonious, but also to avoid traps associated with a misspecified model. Lamoureux and Zhang (2024) show that despite its parsimony this algorithm can suffer from overfitting. They show that estimation risk lives in the portfolio variance, so more risk tolerant investors’ optimal portfolios produce large losses out of sample.
@@ 3.1 Metropolis within Gibbs draws
$p_{\lambda}(\theta|\mathrm{data})\propto e^{\left\{\lambda U(\mathrm{data},\theta)\right\}}\pi(\theta)=\operatorname*{arg\,min}_{q}\left\{-\int\lambda U((\mathrm{data},\theta)q(\theta)d\theta+\mathrm{KL}(q(\theta)||\pi(\theta)\right\}$ (4)
@@ 3.2 $\lambda$ selection
I develop an algorithm to select the optimal value of $\lambda$, $\lambda^{*}$ by analyzing the properties of the posterior covariance matrix of $\theta$, $\Sigma$. As noted above, at low levels of $\lambda$ ($\lambda=500$), the Gibbs posterior fully reflects the prior—the data has little or no effect. At very high levels of $\lambda$ ($\lambda=100,000$), the Gibbs posterior is tightly distributed about the empirical optimal with no regularization or discipline from the prior. Selecting an optimal value for $\lambda$ entails finding an optimal tradeoff between precision (learning from the data) and overfitting (fragility). I use the condition number of $\Sigma$ (the ratio of the largest to the smallest eigenvalue) to characterize overfitting and $\log\det\Sigma$) to measure precision. I characterize the tradeoff with an identification frontier to select the value of $\lambda$, $\lambda^{*}$, where the marginal increase in precision equals the marginal loss from overfitting under the investor’s utility function.
@@
Figure 1 shows the identification frontier on the $\lambda$ grid for the case of log utility conditional on data from the 240 months from January 1977 through December 1996, and $\lambda^{*}$ is 7,500. Figure 2 shows the information from the same data for the power utility function with coefficient of relative risk aversion, $\gamma=6$, and $\lambda^{*}$ is 2,500. In Figure 1 (Figure 2) at $\lambda=500$, $\Sigma$’s condition number is 1.32 (2.96) and the approximated second derivative is -1.65 (-0.40). At $\lambda=10,000$, these values are 6.99 (23.28) and 3.20 (7.73), respectively. Both figures show that at values of $\lambda$ below $\lambda^{*}$, increasing $\lambda$ has a large effect on the log determinant without increasing the condition number dramatically. By contrast, gains in precision from increasing $\lambda$ above $\lambda^{*}$ are overwhelmed by large increases in the condition number.
@@
This shows the sense in which my precision metric is a proxy for information gain (i.e., entropy reduction). The Hessian determines the curvature of the utility surface around the mode and Equation (8) links the Hessian to the posterior covariance matrix. So the posterior covariance’s condition number summarizes the heterogeneity of that curvature along the directions of the six characteristics. It is in this sense that trading off fragility against the information gain avoids anisotropic entropy.
@@ 4.3 Optimal portfolio return properties
Table 3, Panel D and Figures 13 and 14 provide the Gibbs posterior out-of-sample return distributions in the two subperiods for the most risk-averse investor I consider. In these two figures the benchmark is the value-weighted portfolio of sample stocks in em both subperiods. In the first subperiod the PPP’s Sharpe ratio, 1.5 [1.3, 1.7][1.3\,,\,1.7] is significantly higher than the benchmark’s 0.6, and tightly distributed around its mean. While the PPP’s Gibbs posterior certainty equivalent, 1.8 [0.8, 2.2][0.8\,,\,2.2] is also siginificantly higher than the benchmark’s 0.7, it is negatively skewed with a large variance. This reflects the extreme left tail in the top panel of Figure 14 combined with this investor’s strong aversion to large losses. The lower panel in Figure 14 shows that left tail is no longer present in the second subperiod. However, in this second subperiod positive returns of 18% or more are 10 times more likely under the benchmark than the out-of-sample PPP. Unlike the more risk tolerant investors, in this second subperiod, the decision-theoretic PPP has a higher certainty equivalent than the benchmark (0.23 vs. 0.20 basis points per month), however the Gibbs posterior mean certainty equivalent is lower (0.11).
@@ 4.4 Optimal portfolios’ factor exposures
Table 4 also reports the percentage of portfolio variances attributed directly to each factor—ignoring covariance effects. More than half of the portfolio variance is outside the span of the Fama-French factors for the log utility investor in both subperiods. In the first subperiod this portfolio has large of exposures to the value and momentum factor, and almost no exposure to the other four factors including the market factor. I omit the covariance terms, which are material, explaining why the percentge explained sums to greater than one.
@@ 5. Conclusions
In this paper I show how to apply the tools of Bayesian estimation and decision making to Brandt, Santa-Clara, and Valkanov’s (2009) parametric portfolio policy—with no restrictions on the data generating process. The idea of optimizing a policy function directly on the data has tremendous appeal. As Brandt, Santa-Clara, and Valkanov (2009) and Aït-Sahalia and Brandt (2001) highlight doing so can streamline estimation and reduce the risks from model misspecification. This has several advantages over approaches that rely on traditional machine learning tools such as the bootstrap. From a statistical perspective I avoid assumptions about the data generating process that a bootstrap requires. From a decision theoretic perspective I have a coherent approach that updates my beliefs in light of new data. I conform to the likelihood principle in that only sample data characterize the decision problem. My priors are placed on the parameters that link characteristics to portfolio weights–not the parameters of a model of the return generating process. From a practical perspective, I regularize (mitigate the effects of overfitting) in-sample, using the Gibbs posterior covariance matrix. This facilitates decision making, as less data is needed to estimate and regularize. It also enables more powerful inference as more data is available at the testing stage.
I show that this in-sample regularization works and link the Gibbs posterior weighting parameter to the curvature of the loss function theoretically and empirically. I obtain the exact Gibbs posterior for optimal portfolio weights under the utility function. I use these to construct the predictive distribution of out-of-sample portfolio returns. Thus I have a fully coherent uncertainty distribution without making any assumptions about the return generating process. The paper’s substantive findings rely solely on the observed data, and confirm and update Lamoureux and Zhang (2024) that the traditional characteristics lose predictive efficacy around the turn of this century.
@@
As in Lamoureux and Zhang (2024), the parametric portfolios perform very well in the 20th Century. A likely explanation for this is that the strategies were largely infeasible given the technological and regulatory constraints of that era. Since the year 2000, not only has information about the usefulness of characteristic-based investing been widely disseminated (McLean and Pontiff 2016) and investors have learned about the data generating process (Martin and Nagel 2022; Nagel 2021), transactions costs have dropped dramatically and information processing speeds have increased. A final note on the effectiveness of the in-sample regularization introduced in this paper, Sharpe ratios and certainty equivalents are significantly lower than the benchmarks in the second subperiod only for the log utility function. The heightened regularization helped to mitigate the drop in efficacy for the more risk averse utility functions relative to Lamoureux and Zhang’s (2024) machine learning approach.

### 06 Mar 2026 v2

@@ 1. Introduction
where $\mathcal{L}$ is the loss function and $\pi(\cdot)$ is the prior. A key implication of loss-based updating is that the scale parameter $\lambda$ is not pinned down by Bayesian conditioning under a likelihood. In standard Bayesian analysis, the likelihood fixes the unit of fit and therefore fixes the update’s scale. Since here, I replace the likelihood with a utility function, I have to choose a learning-rate (or “temperature”) parameter, $\lambda$, and its calibration "moves from a modeling detail to a structural component of principled analysis" (McAllin and Takanashi 2026, p. 2). Rather than treating $\lambda$ as a nuisance parameter, I treat it as the central regularization device, and solve for the optimal value of $\lambda$, $\lambda^{\ast}$, in sample.
@@
I obtain posteriors from US equities over the period 1955-2024. I confirm Lamoureux and Zhang’s finding that characteristic tilts afford large utility gains prior to 2001 for investors with varying tolerance for risk. These gains vanish in the 21st Century. I find that $\lambda^{*}$ varies meaningfully with risk aversion and departs from the quadratic benchmark in ways that are informative about higher-order moments. The pattern of departure is stable prior to 2001, and changes in ways consistent with structural changes in the data generating process in the 21st Century.
@@
The paper also contributes to the burgeoning literature that represents returns and risks in characteristic space. Like Kelly, Pruitt, and Su’s (2019) Instrumented Principal Components Analysis (IPCA), I shift the investment universe from a set of CUSIPs to a cloud of observations in characteristic space. In contrast to IPCA, which uses this transformation to identify latent statistical factors, I use the geometry of characteristic-managed returns to construct a Gibbs posterior over utility-maximizing portfolios. Kozak, Nagel, and Santosh (2018, 2020) argue that characteristics span the investment opportunity set and estimate the associated low-dimensional structure via shrinkage and low-rank methods. In a complementary way, their work tells us what returns look like in characteristic space and IPCA tells us how risk moves in characteristic space; this paper tells us how to optimally invest in characteristic space under uncertainty.
@@ 3. Bayesian updating
Brandt, Santa-Clara, and Valkanov (2009) stress that an investor can use this algorithm to maximize in-sample utility without taking a stand on the data generating process. By reducing the portfolio problem to that of estimating the $K$-vector $\theta$ of characteristic tilts, thereby avoiding the estimation of moments, or a complex likelihood function, they hope to mitigate the problem of overfitting that plagues traditional portfolio optimization. As described above, as a Bayesian decision maker, I also want to be completely agnostic about the data generating process—not simply to make the estimation more parsimonious, but also to avoid traps associated with a misspecified model. Lamoureux and Zhang (2024) show that despite its parsimony this algorithm can suffer from overfitting. They show that estimation risk lives in the portfolio variance, so more risk tolerant investors’ optimal portfolios produce large losses out of sample.
@@ 3.1 Metropolis within Gibbs draws
$p_{\lambda}(\theta|\mathrm{data})\propto e^{\left\{\lambda U(\mathrm{data},\theta)\right\}}\pi(\theta)=\operatorname*{arg\,min}_{q}\left\{-\int\lambda U(\mathrm{data},\theta)q(\theta)d\theta+\mathrm{KL}(q(\theta)||\pi(\theta))\right\}$ (4)
@@ 3.2 $\lambda$ selection
I develop an algorithm to select the optimal value of $\lambda$, $\lambda^{*}$ by analyzing the properties of the posterior covariance matrix of $\theta$, $\Sigma$. As noted above, at low levels of $\lambda$ ($\lambda=500$), the Gibbs posterior fully reflects the prior—the data has little or no effect. At very high levels of $\lambda$ ($\lambda=100,000$), the Gibbs posterior is tightly distributed about the empirical optimal with no regularization or discipline from the prior. Selecting an optimal value for $\lambda$ entails finding an optimal tradeoff between precision (learning from the data) and overfitting (fragility). I use the condition number of $\Sigma$ (the ratio of the largest to the smallest eigenvalue) to characterize overfitting and $\log\det\Sigma$ to measure precision. I characterize the tradeoff with an identification frontier to select the value of $\lambda$, $\lambda^{*}$, where the marginal increase in precision equals the marginal loss from overfitting under the investor’s utility function.
@@
Mapping this information frontier in $(\kappa,\frac{d^{2}(-logdet\Sigma)}{d\kappa^{2}}$ space isolates the phase transition in the specific investor’s Bayesian updating. KNEEDLE identifies the inflection point, which shows where the marginal gain in learning from the data (signal quality) starts to decline relative to the marginal increase in fragility. In this way the resulting $\lambda^{*}$ is scaled to this specific investor’s utility function.
@@
Figure 1 shows the identification frontier on the $\lambda$ grid for the case of log utility conditional on data from the 240 months from January 1977 through December 1996, and $\lambda^{*}$ is 7,500. Figure 2 shows the information from the same data for the power utility function with coefficient of relative risk aversion, $\gamma=6$, and $\lambda^{*}$ is 2,500. The two information frontiers are explicit functions of each utility function. In Figure 1, for the log utility investor, at $\lambda=500$, $\Sigma$’s condition number is 1.32 and the approximated second derivative is -1.65 (-0.40), and at $\lambda=10,000$, these values are 6.99 (23.28) and 3.20. In Figure 2 for the power utility investor with $\gamma=6$, the condition number is 2.96 when $\lambda=500$ and the approximate second derivative is -0.40; and at $\lambda=10,000$, the condition number is 7.73 and the approximate second derivative is 23.28. Both figures show that at values of $\lambda$ below $\lambda^{*}$, increasing $\lambda$ has a large effect on the log determinant without increasing the condition number dramatically. By contrast, gains in precision from increasing $\lambda$ above $\lambda^{*}$ are overwhelmed by large increases in the condition number.
@@
This shows the sense in which my precision metric is a proxy for information gain (i.e., entropy reduction). The Hessian determines the curvature of the utility surface around the mode and Equation (8) links the Hessian to the posterior covariance matrix. So the posterior covariance’s condition number summarizes the heterogeneity of that curvature along the directions of the six characteristics. It is in this sense that trading off fragility against the information gain avoids anisotropic entropy. In the mean-variance case the Hessian is strictly proportional to $Q$, the covariance matrix of characteristic-managed portfolio returns, and so $\lambda^{*}$ scales proportionately to $\gamma$. However for power utility functions higher order moments affect the Hessian, which motivates the KNEEDLE algorithm to select $\lambda^{*}$ for each investor and each conditioning data set.
@@ 4.3 Optimal portfolio return properties
Table 3, Panel D and Figures 13 and 14 provide the Gibbs posterior out-of-sample return distributions in the two subperiods for the most risk-averse investor I consider. In these two figures the benchmark is the value-weighted portfolio of sample stocks in both subperiods. In the first subperiod the PPP’s Sharpe ratio, 1.5 [1.3, 1.7][1.3\,,\,1.7] is significantly higher than the benchmark’s 0.6, and tightly distributed around its mean. While the PPP’s Gibbs posterior certainty equivalent, 1.8 [0.8, 2.2][0.8\,,\,2.2] is also significantly higher than the benchmark’s 0.7, it is negatively skewed with a large variance. This reflects the extreme left tail in the top panel of Figure 14 combined with this investor’s strong aversion to large losses. The lower panel in Figure 14 shows that left tail is no longer present in the second subperiod. However, in this second subperiod positive returns of 18% or more are 10 times more likely under the benchmark than the out-of-sample PPP. Unlike the more risk tolerant investors, in this second subperiod, the decision-theoretic PPP has a higher certainty equivalent than the benchmark (0.23 vs. 0.20 basis points per month), however the Gibbs posterior mean certainty equivalent is lower (0.11).
@@ 4.4 Optimal portfolios’ factor exposures
The large statistically significant $\alpha$ estimates have two sources. First, two of the characteristics, month-of-the-year return and residual volatility, are not accounted for by the six-factor model. More importantly, the additive, affine nature of standard factor models may be inadequate to analyze a non-linear utility-optimized strategy. The PPP framework exploits complementarities between characteristics, for example the joint distribution of momentum and size, etc. The large $\alpha$ provides evidence that these cross-characteristic synergies were large in the 20th Century.
Table 4 also reports the percentage of portfolio variances attributed directly to each factor—ignoring covariance effects. More than half of the portfolio variance is outside the span of the Fama-French factors for the log utility investor in both subperiods. In the first subperiod this portfolio has large of exposures to the value and momentum factor, and almost no exposure to the other four factors including the market factor. I omit the covariance terms, which are material, explaining why the percentage explained sums to greater than one.
@@ 5. Conclusions
In this paper I show how to apply the tools of Bayesian estimation and decision making to Brandt, Santa-Clara, and Valkanov’s (2009) parametric portfolio policy—with no restrictions on the data generating process. The idea of optimizing a policy function directly on the data, without specifying a likelihood function, has tremendous appeal. As Brandt, Santa-Clara, and Valkanov (2009) and Aït-Sahalia and Brandt (2001) highlight, doing so can streamline estimation and reduce the risks from model misspecification. My Gibbs posterior has several advantages over approaches that rely on traditional machine learning tools such as the bootstrap. From a statistical perspective I avoid assumptions about the data generating process that a bootstrap requires. From a decision theoretic perspective I have a coherent approach that updates my beliefs in light of new data. I conform to the likelihood principle in that only sample data characterize the decision problem. My priors are placed on the parameters that link characteristics to portfolio weights–not the parameters of a model of the return generating process. From a practical perspective, I regularize (mitigate the effects of overfitting) in-sample, using the geometry of the Gibbs posterior covariance matrix. This facilitates decision making, as less data is needed to estimate and regularize. It also enables more powerful inference as more data is available at the testing stage.
@@
I show that this in-sample regularization works and I link the Gibbs posterior weighting parameter to the curvature of the loss function both theoretically and empirically. In the mean-variance case, the product of the optimal weight on the data and the risk aversion coefficient is constant across investors. Higher order moment structure of the Hessian breaks this relationship in the power utility case. I obtain the exact Gibbs posterior for optimal portfolio weights under the utility function. I use these to construct the predictive distribution of out-of-sample portfolio returns. Thus I have a fully coherent uncertainty distribution without a likelihood. The paper’s substantive findings rely solely on the observed data, and confirm and update Lamoureux and Zhang (2024) that the traditional characteristics lose predictive efficacy around the turn of this century. A likely explanation for this is that the strategies were largely infeasible given the technological and regulatory constraints of that era. Since the year 2000, not only has information about the usefulness of characteristic-based investing been widely disseminated (McLean and Pontiff 2016) and investors have learned about the data generating process (Martin and Nagel 2022; Nagel 2021), transactions costs have dropped dramatically and information processing speeds have increased. A final note on the effectiveness of the in-sample regularization introduced in this paper, Sharpe ratios and certainty equivalents are significantly lower than the benchmarks in the second subperiod only for the log utility function. The heightened regularization helped to mitigate the drop in efficacy for the more risk averse utility functions relative to Lamoureux and Zhang’s (2024) machine learning approach.